OpenStudy (anonymous):

Find the rate of change between the two points (1, 6.2) and (3, 20) where x is in hours and y is in miles. need the answer asap!

OpenStudy (fifciol):

rate of change between points A and B =\[\frac{ y_B - y_A }{x_B-x_A }\]

OpenStudy (fifciol):

Let point A be (1,4). What is \[x_A\] and\[y_A\]?

OpenStudy (fifciol):

well x_A =1 and y_A=4 Substitute in the equation of rate of change values of coordinates, no matter whether you choose first point as A or not

OpenStudy (fifciol):

Do you know how to do that?

OpenStudy (fifciol):

well it's \[\frac{ 3-1 }{ 20-6.2 }\]

OpenStudy (fifciol):

so its\[\frac{ 2 }{ 13,8 }=0,145\]